How many grams of iron oxide can be produced with 56.3 grams of o2 and 57.8 grams of fe

Basic explanations needed
1
Expert's answer
2020-04-13T02:27:30-0400

The number of the moles of iron is:

"n=\frac{m}{M} = \frac{57.8\text{ g}}{ 55.845\text{ g/mol}} = 1.035 \text{ mol}"

The number of the moles of oxygen is:

"n=\frac{m}{M} = \frac{56.3\text{ g}}{ 32.00\text{ g/mol}} = 1.76 \text{ mol}"


Let's write the reaction equation and balance it:

4Fe + 3O2 "\rightarrow" 2Fe2O3


As you can see, 4 moles of iron react with 3 moles of oxygen. Therefore, the number of the moles of O2 needed to react with all the iron available is: "1.035\cdot3/4 = 0.777" mol. The amount of the oxygen is larger: 1.76 mol, so the oxygen is in excess. The number of the moles of the iron oxide Fe2O3 produced when all iron is consumed is:

"n = 1.035\cdot2/4 = 0.518" mol

Therefore, its mass is:

"m = n\cdot M = 0.518 \text{ mol}\cdot159.69 \text{ g/mol} = 82.6" g

Answer: 82.6 g of iron oxide can be produced with 56.3 grams of O2 and 57.8 grams of Fe.